### v1

# A Comprehensive Survey of Machine Unlearning Techniques for Large Language Models

Abstract
This study investigates the machine unlearning techniques within the context of large language models (LLMs), referred to as LLM unlearning. LLM unlearning offers a principled approach to removing the influence of undesirable data (e.g., sensitive or illegal information) from LLMs, while preserving their overall utility without requiring full retraining. Despite growing research interest, there is no comprehensive survey that systematically organizes existing work and distills key insights; here, we aim to bridge this gap. We begin by introducing the definition and the paradigms of LLM unlearning, followed by a comprehensive taxonomy of existing unlearning studies. Next, we categorize current unlearning approaches, summarizing their strengths and limitations. Additionally, we review evaluation metrics and benchmarks, providing a structured overview of current assessment methodologies. Finally, we outline promising directions for future research, highlighting key challenges and opportunities in the field.
1 Introduction
Figure 1: Overview of LLM Unlearning. LLM unlearning focuses on removing specific data (forget set) while minimizing the impact on related knowledge (retain set) and general world knowledge.
Figure 2: Taxonomy of LLM unlearning.
The widespread adoption of large language models (LLMs) has brought significant challenges, particularly concerning user data privacy, copyright protection, and alignment with societal values. During training, these models can inadvertently memorize sensitive information, such as personally identifiable data or copyrighted materials Li et al. (2024a, b); Zhang et al. (2024b); Yao et al. (2024). In addition to privacy and copyright issues, some training data may embed content that conflicts with contemporary social norms, such as discriminatory language based on race, ethnicity, etc Li et al. (2025). These biases often manifest as harmful stereotypes, undermining the fairness and inclusivity of AI systems. Addressing these concerns is not only a societal imperative but also a regulatory requirement under privacy laws such as the General Data Protection Regulation (GDPR) Council of the European Union (2016) and the EU Artificial Intelligence Act (EU AI Act) Council of the European Union (2024). These laws mandate the “right to be forgotten” and require mechanisms to delete specific data upon request.
To address these challenges, the field of LLM unlearning has emerged, focusing on removing specific information or behaviors from models while preserving their overall performance. However, LLM unlearning faces significant technical challenges. One of the most pressing issues is the prohibitively high cost of retraining. Traditionally, addressing harmful or unwanted data required retraining the model from scratch after excluding problematic data from the training set Jang et al. (2023). This is impractical for LLMs due to the immense time and computational resources required Li et al. (2024a). Moreover, the frequent unlearning requests that arise in deployed models highlight the need for more efficient unlearning techniques. The complexity of LLMs, with their millions or even hundreds of billions parameters, further complicates the task of removing specific information without causing unintended side effects, such as performance degradation or catastrophic forgetting Zhang et al. (2024a).
Despite recent advances in LLM unlearning Eldan et al. (2024); Maini et al. (2024), there is no comprehensive review of existing approaches and evaluation measures. Here we aim to bridge this gap. In particular, we offer a thorough overview of the field, including various unlearning and evaluation methods, and we make the following contributions:
  • We formalize the LLM unlearning paradigms and propose a comprehensive taxonomy to categorize the existing approaches. This taxonomy not only offers a structured understanding of the research landscape, but also helps researchers identify their areas of interest.
  • We systematically review the existing methods, analyzing their features, strengths, and weaknesses. We further discuss the existing evaluation measures and benchmarks, highlighting the challenges when balancing utility preservation and forgetting quality.
  • We discuss future research opportunities for LLM unlearning, including extending techniques to multimodal models and addressing real-world unlearning requests. These avenues aim to advance the field and address emerging challenges.
2 Preliminaries and Taxonomy
2.1 Problem Definition
The objective of LLM unlearning is to selectively remove the influence of specific information while maintaining the model’s overall utility for other tasks. The optimization objective of the model parameters $\theta$ can be expressed as follows:
$\min_{\theta}\mathcal{L}(\theta)=\min_{\theta}\{-\mathcal{L}_{f}(\theta)+\lambda\mathcal{L}_{r}(\theta)\}$ (1)
Here, the forget loss $\mathcal{L}_{f}(\theta)$ quantifies the model prediction error on the forget set $\mathcal{D}_{f}$, while the retain loss $\mathcal{L}_{r}(\theta)$ ensures the preservation of the model’s utility on the retain set $\mathcal{D}_{r}$. The regularization parameter $\lambda\geq 0$ controls the trade-off between effectively forgetting undesired information and preserving the model’s utility.
2.2 LLM Unlearning Paradigms
LLM unlearning follows two main paradigms. The fine-tuning-then-unlearning paradigm focuses on eliminating knowledge introduced during fine-tuning. As illustrated in Figure 1, this paradigm typically leverages synthetic data (e.g., TOFU Maini et al. (2024) and FIUBench Ma et al. (2024)) and partitions a task-specific dataset into a forget set $\mathcal{D}_{f}$ and a retain set $\mathcal{D}_{r}$ to highlight the unlearning precision. The original model $\theta_{ori}$ is first obtained by fine-tuning a pretrained model $\theta_{pre}$ on the task-specific data to encode the target knowledge. Unlearning techniques are then applied to reduce the model’s reliance on the forget set, resulting in the unlearned model $\theta_{unl}$. The direct-unlearning paradigm focuses on eliminating knowledge acquired during the pretraining stage of $\theta_{ori}$, assuming the target knowledge originates from multiple points within the pretraining dataset. The forget and retain sets are typically sampled from pretraining corpora Yao et al. (2024) or publicly available datasets such as Wiki Jin et al. (2024). This paradigm is also extensively applied in safety alignment tasks, where it aims to eradicate hazardous knowledge and to mitigate risks such as misuse or jailbreak attacks Li et al. (2024a); Zhang et al. (2024b).
Substitutes for Knowledge to be Forgotten
Existing unlearning paradigms primarily specify the knowledge to be forgotten, while leaving the substitutes undefined. In this case, refusal responses, such as “I don’t know the answer,” are commonly used as substitutes Maini et al. (2024); Ma et al. (2024)Yao et al. [2024] explored two relabeling strategies. The first strategy uses random labels generated by a randomly initialized model, while the second involves sampling adversarial examples from the model. For each token $w_{t}$ to be unlearned, an adversarial sample $a_{t}$ is generated to maximize the model’s confusion while staying close to $w_{t}$. Eldan et al. [2024] proposed a fine-grained relabeling approach. Their method first identifies tokens whose logit probabilities increase after fine-tuning on the forget set. To facilitate unlearning, a dictionary generated by GPT-4 replaces specific entities with contextually appropriate generic alternatives (e.g., replacing “Hogwarts” with “Mystic Academy”), ensuring more natural and context-aware substitutions.
2.3 Taxonomy
We present a comprehensive taxonomy of LLM unlearning, as illustrated in Figure 2, outlining existing research from the perspectives of methods, evaluation measures, and benchmarks. Existing methods can be categorized into four types: direct fine-tuning, localized parameter modification, leveraging auxiliary models, and input/output-based unlearning. Forgetting quality and utility preservation are critical measures for evaluating unlearning algorithms, particularly given recent discussions on whether knowledge is robustly forgotten or remains susceptible to adversarial recovery. This is often assessed through input-based or logit-based evaluation, as well as model intervention techniques. Additionally, we review commonly used unimodal and multimodal benchmarks.
3 LLM Unlearning Methods
3.1 Direct Fine-Tuning
Relabeling-Based Fine-Tuning first replaces the original responses by generic or neutral substitutes, as described earlier. Then, the model is fine-tuned on the relabeled data to reduce the influence of undesired training information Jin et al. (2024); Eldan et al. (2024).
Reinforcement Learning
The Quark method Lu et al. (2022) is a pioneering approach to applying reinforcement learning for LLM unlearning. It uses a reward model and Proximal Policy Optimization (PPO) Schulman et al. (2017) to reduce undesirable behaviors such as toxicity, repetition, and unwanted sentiment. The reward model assesses the output quality using task-specific measures of toxicity and sentiment. Quark alternates between collecting samples, sorting them into reward-based quantiles labeled with reward tokens, and applying language modeling loss conditioned on these tokens, with a KL-divergence penalty to stay close to the original model. Kassem et al. [2023] proposed DeMem, which leverages a negative similarity reward. This approach trains the LLMs to develop a paraphrasing policy on the forget dataset, generating dissimilar tokens that minimize memorization while preserving semantic coherence.
Gradient Ascent (GA)
Jang et al. [2023] performed gradient ascent (GA) on the next-token prediction loss over the forget set, which is mathematically equivalent to performing gradient descent on the negative prediction loss:
$\small\mathcal{L}_{GA}(\theta)=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{f}}\left[-\log\left(p(y\mid x;\theta)\right)\right].$ (2)
However, the GA-based approach often results in sizable model degradation, characterized by the generation of uniform and low-quality outputs. To mitigate this and to preserve model utility,  Liu et al. [2022] introduced the gradient descent (GD) loss on the retain set $\mathcal{D}_{r}$ as a regularization:
$\small\mathcal{L}_{GD}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{r}}\left[-\log\left(p(y\mid x;\theta)\right)\right].$ (3)
Yao et al. [2024] incorporated the Kullback-Leibler (KL) divergence term as the regularization, which captures the difference between the predictions of the original model and the fine-tuned model on the retain set:
$\small\textstyle\mathcal{L}_{KL}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{r}}\left[D_{KL}\left(p(y\mid x;\theta_{ori})\,\|\,p(y\mid x;\theta)\right)\right].$ (4)
These regularization techniques can also be applied in other unlearning methods to preserve the model’s utility.
Preference Optimization was first designed to align model behavior to human-defined preferences. Specifically, it leverages pairwise comparisons or ranking data to guide the model toward producing outputs that best match desired preferences. Given a preference dataset $\mathcal{D}_{p}=\left\{(x_{i},y_{i,w},y_{i,l})\right\}_{i\in[n]}$, where $y_{i,w}$ and $y_{i,l}$ represent responses to input $x_{i}$, the preference $y_{i,w}>y_{i,l}$ is derived from human comparisons. Direct preference optimization (DPO) Rafailov et al. (2023) minimizes the following objective function:
$\mathcal{L}_{\text{DPO},\beta}(\theta)=$ $-\frac{1}{\beta}\mathbb{E}_{\mathcal{D}_{p}}\left[\log\sigma\left(\beta\log\frac{p(y_{w}\mid x;\theta)}{p(y_{w}\mid x;\theta_{ori})}\right.\right.$ (5)
$-\left.\left.\beta\log\frac{p(y_{l}\mid x;\theta)}{p(y_{l}\mid x;\theta_{ori})}\right)\right],$
where $\sigma$ is the sigmoid function and $\beta$ is the inverse temperature controlling the preference strength. Maini et al. [2024] pioneered the application of DPO to unlearning by framing the forget set as a preference set. The original responses are denoted as $y_{l}$, while refusal responses, such as “I do not know the answer,” are designated as $y_{w}$. This formulation guides the unlearning process by aligning the model’s behavior with the preferred alternative responses. Inspired by this idea, Zhang et al. [2024a] proposed negative preference optimization (NPO), a DPO variant that uses only negative responses from the forget set, disregarding $y_{w}$ in Eq. (5):
$\small\textstyle\mathcal{L}_{\text{DPO},\beta}(\theta)=-\frac{2}{\beta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{f}}\left[\log\sigma\left(-\beta\log\frac{p(y\mid x;\theta)}{p(y\mid x;\theta_{ori})}\right)\right].$ (6)
Zhang et al. [2024a] further theoretically showed that NPO converges to GA as $\beta\to 0$ and the speed toward collapse using NPO is exponentially slower than GA.
3.2 Localized Parameter Modification
Representation Engineering
RMU Li et al. (2024a) focuses on unlearning hazardous knowledge in LLMs by fine-tuning the lower layer $l$ to redirect internal representations of token $t$ in the forget set toward a fixed-noise vector $\mathbf{u}$:
$\small\mathcal{L}_{\text{forget}}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{f}}\frac{1}{L_{x}}\sum_{t\in x}\|\mathcal{H}^{(l)}_{\theta}(t)-c\cdot\mathbf{u}\|_{2}^{2},$ (7)
where $L_{x}$ is the number of tokens in $x$ and $c$ is some hyperparameter that controls activation scaling, $\mathcal{H}^{(l)}_{\theta}(t)$ denotes the internal activations of token $t$ at layer $l$. Simultaneously, it ensures that preserved knowledge remains consistent with the original model by aligning its representations, denoted as
$\small\mathcal{L}_{\text{retain}}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{r}}\frac{1}{L_{x}}\sum_{t\in x}\|\mathcal{H}^{(l)}_{\theta}(t)-\mathcal{H}^{(l)}_{\theta_{\text{ori}}}(t)\|_{2}^{2}.$ (8)
Huu-Tien et al. [2024] proposed adaptive RMU, which dynamically scales the random unit vector $\mathbf{u}$ with an adaptive coefficient $\beta\|\mathcal{H}_{\theta_{\text{ori}}}^{(l)}(x)\|_{2}^{2}$ for improved unlearning across layers, unlike RMU’s fixed-scaling coefficient $c$.
Locate-Then-Unlearn Methods
This method focuses on identifying and localizing key model components (e.g., layers or neurons) that are critical for unlearning. DEPN Wu et al. (2023) leverages privacy attribution via gradient integration to identify privacy-sensitive neurons. It quantifies each neuron’s contribution to privacy leakage by efficiently approximating the integral of the gradient changes using a limited number of steps (e.g., $m=20$). Specifically, the privacy attribution score $\text{Att}(w_{k}^{l})$ for a neuron $w_{k}^{l}$ at layer $l$ is computed using the following cumulative gradient integration:
[…]
where $\beta_{k}^{l}$ is the activation value of the neuron, $\alpha_{k}^{l}$ represents the modified activation value, and $p(y|x)$ is the conditional probability of the model predicting the private information.
WAGLE Jia et al. (2024) uses bi-level optimization to examine the interaction between weight adjustment and unlearning efficacy. By leveraging weight attribution, it quantifies the relationship between weight influence and the impact of forgotten or retained data on LLM outputs. The unlearning sensitivity score for weight perturbation is obtained from the forget loss $S_{i}=\mathcal{L}_{f}(\epsilon\odot\theta(\epsilon))-\mathcal{L}_{f}(\theta(1))$, where $\epsilon\odot\theta(\epsilon)$ is the weight-adjusted model, $\epsilon$ represents weight modifications, and $\epsilon=1$ indicates no interference. The weights $\theta(\epsilon)$ minimize the retain loss $\mathcal{L}_{r}$ to preserve utility. WAGLE uses a diagonal Hessian approximation for computational efficiency and accuracy, with the sensitivity score expressed as
$S_{i}\propto[\theta]_{i}[\nabla\mathcal{L}_{f}(\theta)]_{i}-\frac{1}{\gamma}[\nabla\mathcal{L}_{r}(\theta)]_{i}[\nabla\mathcal{L}_{f}(\theta)]_{i},$ (10)
where $[\theta]_{i}$ is the $i$-th weight, $[\nabla\mathcal{L}_{f}(\theta)]_{i}$ and $[\nabla\mathcal{L}_{r}(\theta)]_{i}$ are the gradients of the forget and the retain losses for the $i$-th weight, respectively, and $\gamma$ is the Hessian parameter.
Beyond DEPN and WAGLE, Guo et al. [2024] introduced a mechanistic unlearning framework that combines fact lookup localization, using logistic regression probes or path patching to assess causal importance, with localized fine-tuning. Similarly, Needle Hong et al. (2024) identifies and disrupts concept vectors in MLP layers encoding specific knowledge using vocabulary projections and causal tests.
Table 1: A summary of existing LLM unlearning studies. C-Utility, C-Forget, C-Efficiency denote the challenges to solve in the perspective of utility preservation, forget quality, and efficiency, respectively.
Methodology Features C-Utility C-Forget C-Efficiency Related Work Application
Relabeling-Based Fine-Tuning Using refusal string or fine-grained alternatives Eldan et al. (2024) Privacy & Copyright
Gradient Ascent Maximizing cross-entropy loss w.r.t target sequence Jang et al. (2023) Privacy & Copyright
Yao et al. (2024) Privacy & Copyright
Preference Optimization Better utility preservation Maini et al. (2024) Privacy & Copyright
Zhang et al. (2024a) Privacy & Copyright
Zhang et al. (2024b) AI Alignment
Reinforcement Learning Need for an accurate reward model Kassem et al. (2023) Privacy & Copyright
Lu et al. (2022) AI Alignment
Representation Engineering Modify localized weights to update internal representation Li et al. (2024a) AI Alignment
Huu-Tien et al. (2024) AI Alignment
Locate-then- Unlearn First locate the parameters corresponding to the target knowledge, then modify them. Wu et al. (2023) Privacy & Copyright
Jia et al. (2024) Privacy & Copyright
Guo et al. (2024) Privacy & Copyright
Hong et al. (2024) Privacy & Copyright
Task Vector Utilize the model weight difference before and after fine-tuning Ilharco et al. (2023) AI Alignment
Liu et al. (2024c) AI Alignment
Gao et al. (2024) AI Alignment
Contrastive Decoding Intervention during the decoding process Ji et al. (2024) AI Alignment
Knowledge Distillation Adjusting logits to guide the student model Wang et al. (2024) Privacy & Copyright
Dong et al. (2024) Privacy & Copyright
Input/Output- Based Modify inputs or outputs instead of model weights. Pawelczyk et al. (2025) Privacy & Copyright
Liu et al. (2024a) Privacy & Copyright
Thaker et al. (2024) Privacy & Copyright
3.3 Leveraging Auxiliary Models
These methods typically fine-tune an assistant model $\theta_{a}$ to replicate knowledge from $\mathcal{D}_{f}$. Its outputs are then used to adjust the original model’s responses, mitigating the influence of $\mathcal{D}_{f}$ through the auxiliary model’s weights or logits.
Contrastive Decoding
ULD Ji et al. (2024) uses an auxiliary model trained on the forget set to guide the unlearning process during decoding. It claims that the auxiliary LLM should exclusively capture the unique knowledge within the forget set while preventing the retention of any information meant to be preserved. Ideally, this results in a uniform distribution over the retain set. The optimization objective of the auxiliary model is formulated as the inverse of Eq. (1):
$\small\min_{\theta_{a}}\mathcal{L}(\theta_{a})=\min_{\theta_{a}}\{\mathcal{L}_{f}(\theta_{a})-\beta\mathcal{L}_{r}(\theta_{a})\},$ (11)
The retain loss $\mathcal{L}_{r}(\theta_{a})$ is specifically formulated as the cross-entropy with respect to the uniform distribution. To enhance the efficiency in the auxiliary model’s implementation, the first $k$ transformer layers of the original LLM are reused, along with the language model head, to map hidden representations to output logits across the entire vocabulary.
Knowledge Distillation
uses a specialized unlearning teacher model to guide the unlearning process, providing signals to the student model to adjust the logits and to selectively forget specific information. RKLD Wang et al. (2024) first identifies tokens requiring unlearning by detecting such with consistently increased logit values after fine-tuning. The formula for the unlearning teacher model is as follows:
$\small l(y|x;\theta_{ori})-\alpha\cdot{ReLU}(l(y|x;\theta_{a})-l(y|x;\theta_{ori}))$ (12)
where $l(y|x;\theta_{ori})$ and $l(y|x;\theta_{a})$ represent the logits of the original and the auxiliary model, respectively, and $\alpha$ is a hyperparameter that controls the forgetting strength, respectively. This strategy offers more precise guidance for intentional forgetting while safeguarding other information. Moreover, Dong et al. [2024] introduced a self-distillation method that assumes tokens like named entities or nouns contain sensitive information requiring unlearning. To identify these key tokens, they used a syntactic analysis tool for extraction. The teacher model’s logits were derived by subtracting hyperparameter-controlled one-hot vectors for the key tokens from the logits of the original model.
Task Vector
Task vectors, defined as $\tau=\theta_{a}-\theta_{ori}$, steer the model behavior by editing the weight space, thus enabling operations such as negation and addition for applications such as unlearning and multi-task learning. Negating task vectors effectively suppress behaviors such as mitigating toxic language generation Ilharco et al. (2023). Liu et al. [2024c] enhanced fine-tuning with modules targeting harmful knowledge: guided distortion, random disassociation, and preservation divergence. Ethos Gao et al. (2024) distinguishes general and undesired knowledge by projecting task vectors onto principal components. By negating only undesired components, Ethos minimizes collateral damage to model utility, thus achieving unlearning.
3.4 Input/Output-Based Unlearning
Input/output-based unlearning methods offer flexibility by using prompt engineering and post-processing without modifying the model weights or the architecture.  Liu et al. [2024a] proposed training a prompt classifier to identify prompts within the scope of unlearning and efficiently corrupting them in the embedding space using zeroth-order optimization. Thaker et al. [2024] showed that simple guardrails like prompting and input/output filtering can effectively support unlearning independently or alongside fine-tuning.  Pawelczyk et al. [2025] proposed in-context unlearning by constructing tailored prompts, where the labels of the data to be forgotten are flipped.
Brief Summary
Table 3.2 summarizes the existing unlearning methodologies, their features, representative work, and applications, highlighting challenges in forget quality, utility preservation, and efficiency based on their principles and empirical findings. Methods such as gradient ascent, preference optimization, and reinforcement learning often struggle with utility preservation due to optimization conflicts (e.g., maximizing cross-entropy loss), requiring regularization to mitigate the performance loss. Relabeling-based fine-tuning and knowledge distillation have been reported to suffer from knowledge relearning. This issue arises when models are fine-tuned using small amounts of related or even unrelated samples, or through in-context learning Lynch et al. (2024). Representation engineering methods have been reported to be more susceptible to adversarial attacks, failing to robustly erase knowledge Łucki et al. (2024). Contrastive decoding and input/output-based unlearning are controversial because they do not truly remove knowledge from the models. Task vectors, limited by imprecision in localizing knowledge, are mainly used in AI alignment. From an efficiency standpoint, reinforcement learning and locate-then-unlearn approaches involve higher algorithmic complexity. Reinforcement learning requires training high-quality models to generate reward signals, while locate-then-unlearn methods rely on precise knowledge localization. Moreover, contrastive decoding and input/output-based methods increase the computational overhead during inference, which leads to slower generation.
4 Benchmarks
This section provides a detailed description of the commonly used benchmarks, addressing areas such as copyright, privacy, and AI alignment. Notably, WMDP and RWKU are designed for the direct-unlearning, while other benchmarks are used within the fine-tuning-then-unlearning paradigm.
4.1 Unimodal Benchmarks
Who is Harry Potter (WHP) Eldan et al. (2024) evaluates unlearning of Harry Potter-related information using a dataset combining the original books (2.1M tokens) with synthetic content (1M tokens). Unlearning effectiveness is assessed through 300 Harry Potter-related questions scored with a GPT-4-based evaluation. It explores the potential of unlearning in copyright protection.
TOFU Maini et al. (2024) includes 200 synthetic author profiles with 20 question-answer examples each, ensuring no overlap with existing training data. In addition to offering three forget-retain splits (90-10, 95-5, and 99-1), the benchmark includes 100 real-world author profiles and 117 world facts, providing a comprehensive evaluation of model utility after unlearning.
WMDP Li et al. (2024a) comprises 3,668 multiple-choice questions on biosecurity, cybersecurity, and chemical security, curated by experts to evaluate hazardous knowledge while excluding sensitive information. It serves as a benchmark for assessing LLMs’ hazardous knowledge and unlearning methods for AI alignment.
RWKU Jin et al. (2024) includes 200 unlearning targets focused on well-known public figures, with a total of 13,131 multi-level forget probes. These probes are divided into 3,268 fill-in-the-blank probes, 2,879 question-answer probes, and 6,984 adversarial attack probes.
4.2 Multimodal Benchmarks
FIUBench Ma et al. (2024) comprises 400 synthetic faces paired with fictitious private data such as personal backgrounds, health records, criminal histories, phone numbers, occupations and incomes are randomly assigned. GPT-4o generates 20 question-answer pairs for each profile.
MLLMU-Bench Liu et al. (2024b) contains 500 fictitious profiles and 153 celebrity profiles, each with 14+ question-answer pairs evaluated in multimodal (image+text) and unimodal (text-only) settings. It features 20.7k questions, with fictitious profiles generated by GPT-4o and real celebrity profiles reviewed by experts. The test set includes 3.5k paraphrased questions and 500 modified images with pose variations. An additional utility set uses celebrity profiles.
CLEAR Dontsov et al. (2024) focuses on person unlearning, characterizing unlearning across textual and visual modalities. It generates consistent images through a comprehensive strategy and links them to the corresponding author-related questions from TOFU. It includes a total of 200 fictitious individuals linked with 3.7k visual question-answer pairs and 4k textual question-answer pairs, enabling a thorough evaluation of unimodal and multi-modal unlearning techniques.
5 Evaluation Measures
This section outlines common evaluation measures used to access both utility preservation and forgetting quality.
5.1 Utility Preservation Evaluation
We categorize the methods for evaluating model utility into general utility and localized utility. General utility evaluates the model’s overall knowledge, including downstream task performance and related metrics. In contrast, localized utility examines fine-grained changes, such as output or logit probability similarity, in the model’s output before and after unlearning on task-specific data.
5.1.1 General Utility
Accuracy on General Benchmarks
Unlearned models must retain their original world knowledge without degradation during the unlearning process. It is common to use additional benchmarks Joshi et al. (2017); Fu et al. (2023) to evaluate the utility preservation.
GPT-Eval
LLM-as-a-Judge has been widely adopted across various evaluation scenarios due to its flexibility and ability to assess multiple dimensions. It provides semantic insights that traditional measures might overlook. For instance, Ma et al. [2024] leverage GPT-4o Mini to evaluate unlearning algorithms based on correctness, helpfulness, and relevance, assigning an overall score on a scale from 0 to 1.
5.1.2 Localized Utility
ROUGE
is a measure for evaluating word-level alignment between a model’s output and the ground truth answer. The ROUGE-L recall score is used to quantify the overlap between the generated output and the reference answer Maini et al. (2024); Yuan et al. (2024).
Probability
Maini et al. (2024) evaluates the model’s ability to predict the correct answer for a given question by computing the normalized conditional probability of the answer, defined as $P(y|x)=\frac{1}{T}\sum_{t=1}^{T}p(y_{t}|y_{<t}\circ x;\theta_{unl})$.
Truth Ratio
evaluates the relative likelihood of a correct answer compared to an incorrect one for a given question Maini et al. (2024). It is defined as the ratio of the average normalized conditional probability of perturbed incorrect answers $\bar{y}$ from the $\mathcal{A}$, to that of paraphrased correct answer $\tilde{y}$ of the ground truth $y$. Ideally, when a model lacks the relevant knowledge, it should assign similar prediction probabilities to both correct and incorrect answers. Formally, the truth ratio on the forget set is defined as $R_{truth}(y)=\frac{\frac{1}{\mathcal{A}}\sum_{\bar{y}\in\mathcal{A}}p(\bar{y}|x;\theta_{unl})^{1/|\bar{y}|}}{p(\tilde{y}|x;\theta_{unl})^{1/|\tilde{y}|}}$, where $\bar{y}$ and $\tilde{y}$ can be generated by LLMs. It is expressed as $\max(0,1-R_{truth}(y))$ on the retain set and world knowledge datasets.
5.2 Forget Efficacy Evaluation
Utility-based evaluation, which compares the utility changes of a model on the forget set before and after unlearning, is the most commonly assessment method. However, recent studies Lynch et al. (2024); Ma et al. (2024) have highlighted the critical need to differentiate between real forgetting and mere suppression, where the suppressed knowledge can often be recovered through adversarial techniques. This section reviews existing methods for comprehensive forgetting quality evaluation beyond utility-based evaluation.
5.2.1 Input-Based Evaluation
Paraphrasing and other Languages
Using generative models such as GPT, paraphrased questions can be easily generated to test the unlearning robustness Maini et al. (2024). Additionally, inspired by the performance disparity of large models across different languages, Lynch et al. [2024] evaluates the robustness of LLM unlearning by translating inputs into Spanish and Russian.
In-Context Relearning
This method leverages demonstrations to interfere with the model’s generation process, prompting it to answer questions it is supposed to forget rather than refusing them. Lynch et al. [2024] show that providing unlearned models with small amounts of general context related to Harry Potter can effectively resurface suppressed knowledge.
Jailbreak Prompts
Jailbreak methods manipulate LLM behavior using black-box techniques like role-playing and obfuscation or white-box strategies that optimize prefixes and suffixes to maximize affirmative response likelihood. In the context of learning, Lynch et al. [2024] evaluated two jailbreak prompts that successfully access the model’s memory of the target dataset through privilege escalation and role-playing.  Łucki et al. [2024] introduced the improved GCG to bypass the unlearning mechanism by optimizing prompt prefixes and leveraging early token loss.
5.2.2 Logit-Based Evaluation
Kolmogorov-Smirnov Test (KS-Test) is a non-parametric statistical method widely used to compare two probability distributions. In the context of forget quality evaluation, the KS-Test assesses the divergence between the distributions of model outputs (e.g., token probabilities or embeddings) before and after unlearning specific data. By computing the maximum absolute difference between the cumulative distribution functions (CDFs) of the two distributions, the test quantifies the extent to which the influence of the unlearned data has been mitigated. A smaller KS statistic indicates a higher degree of successful unlearning. The test’s non-parametric nature ensures its robustness across diverse datasets and scenarios, offering a reliable evaluation measure Maini et al. (2024); Ma et al. (2024).
Membership Inference Attacks (MIAs) determine whether a specific data point is part of a model’s training data (member) or originates from outside the training set (non-member). Therefore, it serves as a tool to evaluate unlearning efficacy Jin et al. (2024); Yao et al. (2024).  Jin et al. [2024] used various MIA methods to assess the robustness of unlearning and found that many unlearning methods failed under such attacks.
5.2.3 Model Intervention
Relearning through Fine-tuning
Hu et al. [2024] suggested to use benign, publicly available data loosely related to the unlearned information. For example, while specific details about virus design might be unavailable, general knowledge such as “What is avian influenza?” is readily accessible. Their experiments demonstrate that augmenting with such benign data can inadvertently cause LLMs to relearn harmful knowledge, a finding further validated by Łucki et al. (2024).
LLM Quantization
facilitates faster inference and efficient deployment on resource-constrained devices with minimal performance degradation. However, recent studies Zhang et al. (2024c); Łucki et al. (2024) indicated that quantization can inadvertently aid in recovering unlearned knowledge. Zhang et al. [2024c] reported that for unlearning methods constrained by utility requirements, unlearned models retain an average of 21% of the intended forgotten knowledge in full precision, which dramatically increases to 83% after 4-bit quantization.
Activation Intervention
Recent research Arditi et al. (2024) demonstrated that refusal behavior in language models is mediated by a single direction in the activation space. Building upon this, outputs from each transformer block of both original and unlearned models are collected using the forget preference dataset Łucki et al. (2024); Seyitoğlu et al. (2024), and the refusal direction for each layer is computed by analyzing the differences in means. During inference, the refusal direction is removed at each layer. Extensive experiments confirmed the effectiveness of this approach.
6 Future Directions
Theoretical framework
Existing LLM unlearning methods often lack formal guarantees of effectiveness. While locate-then-unlearn approaches Wu et al. (2023); Jia et al. (2024) enhance interpretability, they do not establish a rigorous theoretical foundation. A crucial future direction is to develop a comprehensive framework that formally defines and ensures its effectiveness. This could involve leveraging principles from information theory Jeon et al. (2024) and other theoretical approaches to provide a more principled understanding of LLM unlearning.
Unlearning in Multimodal Models
While numerous multimodal datasets have been introduced for multimodal unlearning, current methods remain largely confined to text-based unlearning approaches Ma et al. (2024); Liu et al. (2024b); Dontsov et al. (2024). Future research should prioritize the development of techniques capable of identifying and isolating modality-specific representations within MLLMs. Additionally, robust evaluation benchmarks are essential for assessing the effectiveness of multimodal unlearning methods in disentangling representations where knowledge is intertwined across both textual and visual inputs.
Real-World Complex Requests
Current unlearning methods primarily focus on removing specific data points from the model, requiring explicit target data points (sequences) to be provided. However, real-world unlearning requests may differ from this assumption. A significant future direction for LLM unlearning lies in addressing more complex requests, such as entity-level unlearning, which aims to remove all knowledge related to a specific entity across diverse contexts and associations. This involves not only forgetting explicit facts, but also erasing implicit or derived knowledge. Choi et al. [2024] introduced datasets to evaluate the effectiveness of algorithms in entity-level unlearning tasks. Looking ahead, even more complex scenarios may emerge, such as removing all information about a specific organization, or erasing entire domains of knowledge such as medical or criminal records.
7 Conclusion
We provided a comprehensive survey of recent advances in large language model (LLM) unlearning. We began by defining the problem and outlining the foundational settings of LLM unlearning. To offer a structured understanding, we proposed a novel taxonomy that categorizes existing research from diverse perspectives. We futher explored the methodologies used to implement unlearning and evaluates the effectiveness of these approaches in achieving the desired forgetting. Finally, we examined the key challenges in the field and identified promising directions for future research, thus offering valuable insights for researchers and practitioners.